# Regret Analysis: a control perspective

### v2

@@ Abstract
Online learning and model reference adaptive control have many interesting intersections. One area where they differ however is in how the algorithms are analyzed and what objective or metric is used to discriminate “good” algorithms from “bad” algorithms. In adaptive control there are usually two objectives: 1) prove that all time varying parameters/states of the system are bounded, and 2) that the instantaneous error between the adaptively controlled system and a reference system converges to zero over time (or at least a compact set). For online learning the performance of algorithms is often characterized by the regret the algorithm incurs. Regret is defined as the cumulative loss (cost) over time from the online algorithm minus the cumulative loss (cost) of the single optimal fixed parameter choice in hindsight. Another significant difference between the two areas of research is with regard to the assumptions made in order to obtain said results. Adaptive control makes assumptions about the input-output properties of the control problem and derives solutions for a fixed error model or optimization task. In the online learning literature results are derived for classes of loss functions (i.e. convex) while a priori assuming that all time varying parameters are bounded, which for many optimization tasks is not unrealistic, but is a non starter in control applications. In this work we discuss these differences in detail through the regret based analysis of gradient descent for convex functions and the control based analysis of a streaming regression problem. We close with a discussion about the newly defined paradigm of online adaptive control and ask the following question “ Are regret optimal control strategies deployable?”
@@ 5 Closing
We close with a question. Are regret optimal control strategies deployable? Any time an exploration signal is active it is unlikely that the controller will be in a deployed state. By deployed state we mean when the controller is needed to actively regulate the system (or track a reference command) in a safety critical or high-performance setting, like an aircraft mid-flight. By minimizing regret in its current paradigm OAC is trying to do two things at once (learn and regulate) which are individually being performed in a suboptimal fashion. If you accept the above claim regarding exploration and a “non-deployed state” then any learning based control system that requires an exploration signal would really only be utilized in an episodic paradigm with a learning phase (identification phase/state) and a control phase (deployed phase/state) [^7] The term “episodic strategy” was suggested to the authors as the appropriate terminology for this style of learning and control but we cant remember who suggested this to us. Please reach out and we will credit you in the acknowledgments. With an episodic paradigm you would want to explore as much as possible during the learning phase and not have a decaying exploration signal. We also want to note that while “online” within the ML community is synonymous with regret analysis, this is not the case within the broader engineering community. If one were hearing the words “online adaptive control” for the first time one might assume that you are referring to a real-time control strategy that adapts while controlling. If you accept our above claim about exploration signals and episodic learning then OAC would not be “adapting” in real-time while it was in a deployed state.

### v3

@@ Abstract
Online learning and model reference adaptive control have many interesting intersections. One area where they differ however is in how the algorithms are analyzed and what objective or metric is used to discriminate “good” algorithms from “bad” algorithms. In adaptive control there are usually two objectives: 1) prove that all time varying parameters/states of the system are bounded, and 2) that the instantaneous error between the adaptively controlled system and a reference system converges to zero over time (or at least a compact set). For online learning the performance of algorithms is often characterized by the regret the algorithm incurs. Regret is defined as the cumulative loss (cost) over time from the online algorithm minus the cumulative loss (cost) of the single optimal fixed parameter choice in hindsight. Another significant difference between the two areas of research is with regard to the assumptions made in order to obtain said results. Adaptive control makes assumptions about the input-output properties of the control problem and derives solutions for a fixed error model or optimization task. In the online learning literature results are derived for classes of loss functions (i.e. convex) while a priori assuming certain signals are bounded. In this work we discuss these differences in detail through the regret based analysis of gradient descent for convex functions and the control based analysis of a streaming regression problem. We close with a discussion about the newly defined paradigm of online adaptive control.
@@ 5 Closing
We hope this note was helpful in familiarizing the control community with some of the technical nuances of regret analysis.